to what amount will $5100 invested for 8 years at 10% compounded annually accumulate?
Answer by stanbon(75887)

to what amount will $5100 invested for 8 years at 10% compounded annually accumulate?
-------
A(t) = P(1+(r/n))^(nt)
------
A(8) = 5100(1+(0.1/1))^(1*8)
-----
A(8) = 5100*1.1^8 = $10932.30
==================
Cheers,
Stan H.